OpenStudy (anonymous):

Help solving please: 7+x-15=-2 2/3

OpenStudy (anonymous):

the right side is it \[-2\frac{ 2 }{ 3 }\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

well just add 7 and -15 first

OpenStudy (anonymous):

I got -8

OpenStudy (anonymous):

ok so the left side would be x-8

OpenStudy (anonymous):

then solve for x using \[x-8=-\frac{ 8 }{ 3 }\]

ganeshie8 (ganeshie8):

@jewjewbri do you knw how to add mixed fractions ?

OpenStudy (anonymous):

No I don't

OpenStudy (anonymous):

well multiply everything by 3 so u take out the fraction

OpenStudy (anonymous):

How did you get the 8 over the 3?

OpenStudy (anonymous):

when i converted 2 (2/3)

OpenStudy (anonymous):

how did you convert it?

OpenStudy (anonymous):

2 is just basically 6/3 so then u just add 2/3 to that and get 8/3

OpenStudy (anonymous):

what about the big 2? I'm lost!

ganeshie8 (ganeshie8):

you can see the mixed fraction like this : \(2\frac{2}{3} = 2 + \frac{2}{3}\)

OpenStudy (anonymous):

oh ok..now would the entire answer be 6 1/3 ?

ganeshie8 (ganeshie8):

nope looks something went wrong in your calculation. could you show ur steps...

OpenStudy (anonymous):

Honestly...I just guessed cause thats an answer. idk how to do it

ganeshie8 (ganeshie8):

okay, see if below method makes sense to you: \(x-8 = -2 \frac{2}{3}\) \(x-8 = -(2 +\frac{2}{3})\) \(x-8 = -2 -\frac{2}{3}\) \(x = -2 -\frac{2}{3} + 8\) \(x = -2 -\frac{2}{3} + 8\) \(x = 6 -\frac{2}{3} \)